Q. Rao Corporation has the following balance sheet. How much net operating working capital does the firm have?

Cash

$   10

   Accounts payable

$  20

Short-term investments

                      30

   Accruals

20

Accounts receivable

50

   Notes payable

    50

Inventory

    40

   Current liabilities

$  90

   Current assets

$130

    Long-term debt

60

Net fixed assets

  100

   Common equity

30

   Retained earnings

    50

Total assets

$230

    Total liab. & equity

$230
Answer #1

Net operating working capital = Operating Current Assets -  Operating Current Liabilities

Operating Current Assets are the short-term assets used to support the operations of a business. It generally includes cash, accounts receivable, and inventory

Operating Current Assets = cash + accounts receivable + inventory

= Current assets - Short-term investments

= 130 - 30

= 100

Operating Current Liabilities = Accounts payable + Accruals

= 20 + 20

= 40

Net operating working capital = Operating Current Assets -  Operating Current Liabilities

= 100 - 40

= $60